# DPCL-Diff: The Temporal Knowledge Graph Reasoning based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning

### v1

@@ Abstract
DPCL-Diff: The Temporal Knowledge Graph Reasoning based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning
Yukun Cao1, Lisheng Wang2, Luobing Huang3thanks: 123The authors are affiliated with the School of Computer Science and Technology, Shanghai University of Electric Power (SUEP), Shanghai 201306, China thanks: 1First author (email: marilyn_cao@163.com)thanks: 2Corresponding author (email: 2207992575@mail.shiep.edu.cn)
1 Introduction
@@
Knowledge graphs (KGs) store a large amount of information about the natural world and have shown great success in many downstream applications, such as natural language processing Hu et al. (2022); Wang et al. (2020), recommendation system Xuan et al. (2023); Yu et al. (2022), and information retrieval Zhou et al. (2022). Traditional KGs typically only contain static snapshots of facts, integrating facts (also known as events) in the form of static relational triples $(s,r,o)$, where $s$ and $o$ represent the subject and object entities, respectively, and $r$ represents the relationship type. However, in the real world, knowledge evolves and constantly exhibits complex temporal dynamics Wang et al. (2017); Yin and Cui (2016), which has inspired the construction and application of Temporal Knowledge Graph (TKG). TKG extends the previous static relational triples $(s,r,o)$ to quaternions $(s,r,o,t)$ with timestamps. Thus, the TKG consists of multiple snapshots where facts in the same snapshot appear simultaneously.
@@
Existing studies have identified two main types of new events in TKG: periodic events that will recur Liang et al. (2023); Li et al. (2021), and events that have not occurred before but may occur in the future Xu et al. (2023). For these new events, TKG reasoning offers fresh perspectives and insights for many downstream applications, such as event prediction Zhou et al. (2023), political decision making Zhang and Zhou (2022), dialogue generation Hou et al. (2024), and text generation Li et al. (2022b). These applications have greatly stimulated intense interest in TKG. In this work, we focus on predicting new facts in future time, a task known as graph extrapolation Chen et al. (2023); Li et al. (2021). Our goal is to predict the missing entities in the query $(s,p,?,t)$ for future timestamps $t$ that still need to be observed in the training set.
@@
Most current work models the structural and temporal characteristics of TKG to capture the specific relationships and temporal dependencies between different events for future event prediction.Many studies Liang et al. (2023); Xu et al. (2023); Li et al. (2021)can predict repetitive or periodic events by referring to the known historical events and distinguishing the different effects of periodic and non-periodic events on reasoning tasks through contrastive learning. However, in actual reasoning tasks, some periodic events share the same head entities and relations, differing only in their tail entities. This results in overly similar representations, making it difficult to distinguish them during reasoning. Additionally, In event-based TKG, new events that have never occurred account for about 40% Jäger (2018). Due to the sparse traces of temporal interactions throughout the timeline of these new events, it is impossible to use reasoning methods based on high-frequency periodic events, resulting in poor reasoning performance for these types of events.
@@
To address the above problem, we propose a diffusion model-based dual-domain periodic contrast learning for timing knowledge graph reasoning (DPCL-Diff), which employs a Graph Node Diffusion model (GNDiff) to reason about new events. GNDiff tackles the issue of sparse interaction traces for new events on the timeline by introducing noise into existing correlated sparse events of the new entities. This diffusion mechanism simulates the real-world occurrence of new events and is gradually refined during the denoising process, guided by pre-trained language models (PLMs) that have been trained on extensive corpora and possess a foundational understanding of background knowledge. Consequently, this approach generates high-quality data that closely aligns with the actual distribution. The diffusion model Ho et al. (2020); Song et al. (2020) is a cutting-edge generative framework with demonstrated powerful capabilities in both image Ramesh et al. (2022); Rombach et al. (2022); Saharia et al. (2022)and text generation Ou and Jian (2024); Gong et al. (2022); Strudel et al. (2022). This provides more correlation information for inferring new events and dramatically improves the ability to reason about new events in TKG. For event inference tasks containing a large amount of temporal interaction information, we propose a novel dual-domain periodical contrast learning (DPCL) method. The periodic and non-periodic event entities related to the current event are mapped into the Poincaré and Euclidean spaces for contrastive learning, respectively. The Poincaré space property is utilized to distinguish the similar periodic event entities better.
@@
• GNDiff introduces a novel approach in TKG. GNDiff leverages pre-trained language models (PLMs) to conduct non-Markov decision chain-based diffusion for new events, generating high-quality data and effectively addressing the challenge of sparse interaction traces.
• We propose the DPCL method, where periodic and non-periodic event entities are mapped into the Poincaré space and Euclidean space to better distinguish similar periodic event entities.
@@
2 Related Work
2.1 Diffusion Model
@@
Diffusion models Ho et al. (2020) are latent variable models designed for continuous data, achieving state-of-the-art sample quality in image and audio generation. To handle discrete data, prior work explored text diffusion models in discrete state spaces and defined diffusion processes for such data Li et al. (2024); Lezama et al. (2022); Austin et al. (2021). Li et al. (2022a) focuses on continuous diffusion modelling for text, enabling gradient-based methods to control generation by introducing continuous latent representations. Ou and Jian (2024) combines discrete diffusion models with denoising techniques, significantly improving text generation quality and diversity in unconditional and conditional tasks.
2.2 Temporal Knowledge Graph Reasoning
@@
TKG reasoning can be approached through interpolation or extrapolation Chen et al. (2023); Li et al. (2021). Interpolation addresses missing events between $t_{0}$ to $t_{n}$,while extrapolation predicts events beyond $t_{n}$. Interpolation models like TeAST Li et al. (2023), Re-Temp Wang et al. (2023), and TempCaps Fu et al. (2022) infer missing relationships but cannot predict future events. Extrapolation methods focus on future event prediction. Know-Evolve Trivedi et al. (2017), the first to learn nonlinear evolutionary entity embeddings, struggles with long-term dependencies. CyGNet Zhu et al. (2021) uses replication generation to identify frequent repetitive events, while CENET Xu et al. (2023) builds on this with contrastive learning to identify highly correlated entities. TARGAN Xie et al. (2023) enhances long-term modeling using attentional mechanisms. Other models, including DynaQ Liu et al. (2024a), EvoGNN Pareja et al. (2020), and TempoR Wu et al. (2020) have improved prediction but still face challenges in jointly capturing temporal evolution, semantic dependencies, and scalability.
3 Method
@@
As shown in Fig. 1, for novel events, the DPCL-Diff model leverages GNDiff to inject noise into the sparse correlated events of the new event entity. This approach simulates the real-world mechanism of new events, generating numerous high-quality data samples. For other events, the DPCL method identifies highly correlated entities through comparative learning and better distinguishes similar periodic event entities using Poincaré space properties. In the following few sections, we describe our proposed approach in detail.
3.1 Preliminaries
@@
TKG $G=\{G_{0},G_{1},...,G_{t},...,G_{T}\}$ is a sequence of KGs, each $G_{t}=\{\mathcal{E},\mathcal{R},\mathcal{T}\}$ contains facts that occur at timestamp $t$. Let $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{T}$ denote a finite set of entities, relation types, and timestamps, respectively. Each fact is a quadruple $(s,r,o,t)$ , where $s\in\mathcal{E}$ is the subject entity, $o\in\mathcal{E}$ is the object entity, and $r\in\mathcal{R}$ is the relation at timestamp $t$ between $s$ and $o$ that occurs, the dimension is $d$. ||×$\mathcal{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the embedding of all entities, and its rows represent the embedding vectors of an entity, such as $s$ and $o$. Similarly,||×$\mathcal{R}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the embedding of all relation types. The TKG inference task aims to predict missing objects by answering a query $q=(s,r,?,t)$ using a given history of TKG.
@@
3.2 Graph Node Diffusion Model
In this section, we introduce GNDiff, showcasing the potential of diffusion models for TKG applications. While previous work has effectively explored the diffusion of discrete textual data, applying diffusion models to the discrete graph nodes in TKGs can be similarly effective. Traditional reasoning methods, which rely on historical data to learn implicit associations, need help to reason about new events due to the sparse historical data associated with them. In contrast, GNDiff employs a diffusion model to simulate the event generation mechanism, producing many high-quality samples for new event inference.
GNDiff consists of a forward diffusion process and a reverse diffusion process. For each quadruple $(s,r,o,t)$, after obtaining the vector representations of $s$, $r$, and $o$, these vectors are concatenated into a single vector, which serves as the input to the graph node diffusion model. We then obtain the sample data $x_{0}\sim q(x_{0})$. Since TKGs involve discrete data, each element of $x_{t}$ is a discrete random variable with $K$ categories. For graph node, $K=\left|\mathcal{V}\right|$ represents the number of node types. Representing $x_{t}$ as a stack of heat vectors, the process of adding noise is denoted as shown in Equation 3:
@@ 3.2 Graph Node Diffusion Model
where $\overline{Q}_{t}=Q_{1}Q_{2}\cdots Q_{t}$. Note that $\odot$ denotes elementwise multiplication, and division is performed rowwise. To be combined with pre-trained denoising language models, we incorporate an absorbing state, allowing for the generation of samples during denoising. We use the masking mechanism to simulate the possible inactive state of the node. The transition matrix adjustment process is shown in Equation 6:
$[Q_{t}]_{i,j}=\begin{cases}1&\mathrm{if}\quad i=j=[\mathrm{M}],\\ \beta_{t}&\mathrm{if}\quad j=[\mathrm{M}],\ \mathrm{i}\neq[\mathrm{M}],\\ 1-\beta_{t}&\mathrm{if}\quad i=j\neq[\mathrm{M}],\end{cases}$ (6)
@@
$p_{\theta}({x}_{0:T})=p({x}_{T})\prod_{t=1}^{T}p_{\theta}({x}_{t-1}\mid{x}_{t},t),$ (7)
@@
$\mathbb{E}_{q}[\sum_{t=2}^{T}\mathrm{D}_{\mathrm{KL}}(q({x}_{t-1}|{x}_{t},{x}_{0})\parallel p_{\theta}({x}_{t-1}|{x}_{t},t))]$
@@
3.3 Dual-Domain Periodic Contrastive Learning
Inspired by Xu et al. (2023), we propose a dual-domain periodic contrastive learning method. Unlike the original approach, we use a dual-domain mapping strategy because we observed that specific queries involve multiple similar periodic events that are challenging to distinguish. To address this issue, we map entities from periodic and non-periodic events to Poincaré and Euclidean space, respectively, to identify highly correlated entities within each domain. By leveraging the spatial characteristics of Poincaré space, we effectively differentiate similar periodic entities.
@@ 3.3 Dual-Domain Periodic Contrastive Learning
$\mathbf{S}_{np}^{s,r}(o_{i})=tanh(\mathbf{W}_{np}{s}\oplus r+\mathbf{b}_{np})\mathbf{E}^{T}-\mathbf{Z}_{t}^{s,r}+d_{{E}}(s,o_{i}),$ (12)
Method ICEWS18 ICEWS14
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
ComplEx(2016) 30.09 21.88 34.15 45.96 24.47 16.13 27.49 41.09
R-GCN(2018) 23.19 16.36 25.34 36.48 26.31 18.23 30.43 45.34
ConvE(2018) 36.67 28.51 39.80 50.69 40.73 33.20 43.92 54.35
RE-Net(2019) 42.93 36.19 45.47 55.80 45.71 38.42 49.06 59.12
xERTE(2020) 36.95 30.71 40.38 49.76 32.92 26.44 36.58 46.05
CyGNet(2021) 17.21 30.97 46.85 27.12 27.43 42.63 57.90 37.65
EvoKG(2022) 29.67 12.92 33.08 58.32 18.30 6.30 19.43 39.37
@@
DiffuTKG(2024) 36.72 25.73 57.81 48.51 36.41 72.75
DPCL-Diff 47.02 40.03 49.93 60.66 66.59 62.89 67.79 73.64
Table 1: TKG entity extrapolation results on four datasets. The time-filtered MRR, H@1, H@3, and H@10 metrics are multiplied by 100. The best results are in bold, and the previous best results are underlined.
$\mathcal{L}_{ce}=-\sum_{q}\log\{\frac{exp(S_{p}^{s,r}(o_{i}))}{\sum\limits exp(S_{p}^{s,r}(o_{j}))}+\frac{exp(S_{np}^{s,r}(o_{i}))}{\sum\limits exp(S_{np}^{s,r}(o_{j}))}\},$ (13)
where $o_{i}$ denotes the ground truth object entity for a given query $q$. The goal of $\mathcal{L}_{ce}$ is to distinguish the ground truth from other values by comparing each scalar in $S_{p}^{s,r}$ and $S_{np}^{s,r}$.
$\mathcal{L}_{sup}=\sum_{q\in I}\frac{-1}{\mid P(i)\mid}\sum_{p\in P(i)}\log\frac{\exp(z_{q}\cdot z_{p}/\tau)}{\sum\exp(z_{q}\cdot z_{a}/\tau)}.$ (14)
where $\tau$ is the temperature parameter set to 0.1 in experiments,$I$ is the set of positive samples,$P(i)$ is the set of queries with the same label in the small batch, $z_{q}$ denotes the embedding vector of query $q$, $z_{p}$ represents the embedding vector of queries sharing the same label as $q$, $z_{a}$ denotes the embedding vector of all non-$q$ queries in the small batch.
3.4 Training Details and Inference
@@
$\mathcal{L}_{\mathrm{dpcl}}=\alpha\cdot\mathcal{L}_{\mathrm{diff}}+(1-\alpha)\cdot(\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{sup}}),$ (15)
@@
$\mathbf{P}_{diff}(o_{i}\mid q)=\prod_{t=1}^{T}p_{\theta}(x_{t-1}\mid x_{t}),$ (16)
@@
$\mathbf{P}_{dpcl}(o_{i}\mid q)=\frac{\exp(s(o_{i}\mid q))}{\sum_{o_{j}\in\mathcal{E}}\exp(s(o_{j}\mid q))},$ (17)
@@
4 Experiments
Method WIKI YAGO
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
ComplEx(2016) 47.84 38.15 50.08 58.46 61.29 54.88 62.28 69.42
R-GCN(2018) 37.57 28.15 39.66 —— 41.3 32.56 44.44 ——
ConvE(2018) 47.57 38.76 50.1 63.74 62.32 56.19 63.97 70.44
RE-Net(2019) 51.97 48.01 52.07 53.91 65.16 63.29 65.63 68.08
xERTE(2020) 58.75 58.46 58.85 59.34 58.75 58.46 58.85 60.48
CyGNet(2021) 47.89 66.44 78.70 58.78 17.21 30.97 46.85 27.12
EvoKG(2022) 50.66 12.21 63.84 64.73 55.11 54.37 81.38 92.73
CENET(2023) 66.87 66.77 66.93 66.98 79.64 78.65 79.91 81.42
DHE-TKG(2024) 51.20 57.47 69.25 62.93 71.00 82.72
RLGNet(2024) 64.34 61.03 66.71 69.51 80.17 76.52 83.57 84.96
DPCL-Diff 68.44 68.41 68.45 68.55 84.45 84.23 84.65 84.75
Table 2: Experimental results of temporal link prediction on two public KGs.
4.1 Experimental Setup
Datasets and Evaluation Metrics. To evaluate the effectiveness of the proposed DPCL-Diff, we report the reasoning performance on four public datasets: ICEWS14 Trivedi et al. (2017), ICEWS18 Jin et al. (2019), WIKI Leblay and Chekol (2018), and YAGO Mahdisoltani et al. (2013). Table 3 shows more detailed information about the datasets. We divided all datasets into three groups, roughly partitioning them into train (80%), validation (10%), and test (10%) sets by timestamp. In our experiments, we chose the mean reversed rank (MRR) and the hit rate @1,3,10 Liang et al. (2023) as the metrics.

### v2

# DPCL-Diff: Temporal Knowledge Graph Reasoning Based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning

@@ Abstract
Introduction
@@
Knowledge graphs (KGs) store a large amount of information about the natural world and have shown great success in many downstream applications, such as natural language processing (Hu et al. 2022; Wang et al. 2020), recommendation system (Xuan et al. 2023; Yu et al. 2022), and information retrieval (Zhou et al. 2022). Traditional KGs typically only contain static snapshots of facts, integrating facts (also known as events) in the form of static relational triples $(s,r,o)$, where $s$ and $o$ represent the subject and object entities, respectively, and $r$ represents the relationship type. However, in the real world, knowledge evolves and constantly exhibits complex temporal dynamics (Wang et al. 2017; Yin and Cui 2016), which has inspired the construction and application of Temporal Knowledge Graph (TKG). TKG extends the previous static relational triples $(s,r,o)$ to quaternions $(s,r,o,t)$ with timestamps. Thus, the TKG consists of multiple snapshots where facts in the same snapshot appear simultaneously.
@@
Existing studies have identified two main types of new events in TKG: periodic events that will recur (Liang et al. 2023; Li et al. 2021), and events that have not occurred before but may occur in the future (Xu et al. 2023). For these new events, TKG reasoning offers fresh perspectives and insights for many downstream applications, such as event prediction (Zhou et al. 2023), political decision making (Zhang and Zhou 2022), dialogue generation (Hou et al. 2024), and text generation (Li et al. 2022b). These applications have greatly stimulated intense interest in TKG. In this work, we focus on predicting new facts in future time, a task known as graph extrapolation (Chen et al. 2023; Li et al. 2021). Our goal is to predict the missing entities in the query $(s,p,?,t)$ for future timestamps $t$ that still need to be observed in the training set.
@@
Most current work models the structural and temporal characteristics of TKG to capture the specific relationships and temporal dependencies between different events for future event prediction.Many studies (Liang et al. 2023; Xu et al. 2023; Li et al. 2021)can predict repetitive or periodic events by referring to the known historical events and distinguishing the different effects of periodic and non-periodic events on reasoning tasks through contrastive learning. However, in actual reasoning tasks, some periodic events share the same head entities and relations, differing only in their tail entities. This results in overly similar representations, making it difficult to distinguish them during reasoning. Additionally, In event-based TKG, new events that have never occurred account for about 40% (Jäger 2018). Due to the sparse traces of temporal interactions throughout the timeline of these new events, it is impossible to use reasoning methods based on high-frequency periodic events, resulting in poor reasoning performance for these types of events.
@@
To address the above issue, we propose a diffusion model-based dual-doma in periodic contrast learning (DPCL-Diff) for temporal knowledge graph reasoning (TKG), using a Graph Node Diffusion model (GNDiff) to infer new events. GNDiff solves the sparse interaction traces for new events by adding noise to correlated sparse events of new entities on the timeline. This diffusion mechanism simulates the real-world occurrence of new events, generating high-quality data that closely align with the actual distribution through gradual correction during the denoising process. It also captures the event graph structure, enriching semantic representations and improving the model’s ability to reflect real-world dynamics. The diffusion model (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) is a state-of-the-art generative framework with proven strength in both image (Ramesh et al. 2022; Rombach et al. 2022; Saharia et al. 2022) and text generation (Ou and Jian 2024; Gong et al. 2022; Strudel et al. 2022). This enhances new event inference and reasoning in TKGs. For tasks with large temporal interaction data, we propose a novel Dual-Domain Periodic Contrastive Learning (DPCL) method. Here, periodic and non-periodic event entities are mapped into Poincaré and Euclidean spaces for contrastive learning, respectively. The Poincaré space is used to better distinguish similar periodic entities, capturing entity relationships more effectively than Euclidean space, leading to more accurate link prediction (Balazevic, Allen, and Hospedales 2019; Zeb et al. 2024).
@@
  • GNDiff is introduced in the field of TKG. GNDiff can perform non-Markov decision chain-based graph node diffusion for new events, which generates a large amount of high-quality graph data related to new events.
  • We propose the DPCL method, where periodic and non-periodic event entities are mapped into Poincaré and Euclidean spaces, respectively, to more accurately capture the relationships between periodic event entities and distinguish similar ones.
@@
Related Work
Diffusion Model
@@
Diffusion models (Ho, Jain, and Abbeel 2020) are latent variable models for continuous data, achieving state-of-the-art image and audio generation quality. For discrete data, prior work explored text diffusion in discrete spaces (Li et al. 2024; Lezama et al. 2022; Austin et al. 2021). Li et al. (2022a) enables gradient-based control of text diffusion with continuous latent representations. Ou and Jian (2024) combines discrete diffusion and denoising to improve text generation quality and diversity in both unconditional and conditional tasks.
Temporal Knowledge Graph Reasoning
@@
TKG reasoning involves interpolation and extrapolation (Chen et al. 2023; Li et al. 2021). Interpolation fills missing events within $t_{0}$ to $t_{n}$, while extrapolation predicts events beyond $t_{n}$. Interpolation models like TeAST (Li, Su, and Gao 2023), Re-Temp (Wang, Han, and Poon 2023), and TempCaps (Fu et al. 2022) address missing relationships but cannot predict future events. Extrapolation methods focus on future events: Know-Evolve (Trivedi et al. 2017) captures nonlinear entity evolution but struggles with long-term dependencies, while CyGNet (Zhu et al. 2021) and CENET (Xu et al. 2023) leverage repetitive patterns and contrastive learning. TARGAN (Xie et al. 2023) improves long-term modeling with attention mechanisms. Models like DynaQ (Liu et al. 2024a), EvoGNN (Pareja et al. 2020), and TempoR (Wu et al. 2020) enhance prediction but face challenges in capturing temporal evolution, semantic dependencies, and scalability.
Method
@@
As shown in Fig. 1, for new events, the DPCL-Diff model leverages GNDiff to inject noise into the sparse correlated events of the new event entity. This approach simulates the real-world mechanism of new events, generating numerous high-quality data samples. For other events, the DPCL method identifies highly correlated entities through contrastive learning and more accurately captures the relationships between periodic event entities using the properties of Poincaré space, thereby distinguishing similar periodic entities. In the following sections, we describe our proposed approach in detail.
Preliminaries
@@
TKG $G=\{G_{0},G_{1},...,G_{t},...,G_{T}\}$ is a sequence of KGs, each $G_{t}=\{\mathcal{E},\mathcal{R},\mathcal{T}\}$ contains facts that occur at timestamp $t$. Let $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{T}$ denote a finite set of entities, relation types, and timestamps, respectively. Each fact is a quadruple $(s,r,o,t)$ , where $s\in\mathcal{E}$ is the subject entity, $o\in\mathcal{E}$ is the object entity, and $r\in\mathcal{R}$ is the relation at timestamp $t$ between $s$ and $o$ that occurs, the dimension is $d$. $\textbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the embedding of all entities, and its rows represent the embedding vectors of an entity, such as $s$ and $o$. Similarly, $\textbf{R}\in\mathbb{R}^{|\mathcal{R}|\times d}$ is the embedding of all relation types. The TKG inference task aims to predict missing objects by answering a query $q=(s,r,?,t)$ using a given history of TKG.
@@
Graph Node Diffusion Model
In this section, we introduce GNDiff, demonstrating the effectiveness of diffusion models in TKGs. While diffusion models have been applied to discrete textual data, they are equally effective for discrete graph nodes in TKGs. Traditional methods, reliant on historical data, struggle with new events due to sparse data. In contrast, GNDiff uses a diffusion model to simulate event generation, producing high-quality samples for new event inference.
GNDiff includes a forward and a reverse diffusion process. For each quadruple $(s,r,o,t)$, the vector representations of $s$, $r$, and $o$ are concatenated, forming a single vector as input to the graph node diffusion model. The sample data $x_{0}\sim q(x_{0})$ is then obtained. Since TKGs are discrete, each element of $x_{t}$ is a discrete random variable with $K$ categories, where $K=\left|\mathcal{V}\right|$ is the number of node types. The noise addition process is represented by a stack of heat vectors as shown in Equation 3:
@@
where $\overline{Q}_{t}=Q_{1}Q_{2}\cdots Q_{t}$. Note that $\odot$ denotes elementwise multiplication, and division is performed rowwise. To incorporate the pre-trained language model into the denoising process, we introduce a specific ”inactive” state as the absorbing state, allowing for the generation of samples during denoising. We use the mask mechanism to simulate the possible inactive state of the node. The transition matrix adjustment process is shown in Equation 6:
$[Q_{t}]_{i,j}=\begin{cases}1&\mathrm{if}\quad i=j=[\mathrm{M}],\\ \beta_{t}&\mathrm{if}\quad j=[\mathrm{M}],\ \mathrm{i}\neq[\mathrm{M}],\\ 1-\beta_{t}&\mathrm{if}\quad i=j\neq[\mathrm{M}].\end{cases}$ (6)
@@
$p_{\theta}({x}_{0:T})=p({x}_{T})\prod\nolimits_{t=1}^{T}p_{\theta}({x}_{t-1}\mid{x}_{t},t).$ (7)
@@
$\mathbb{E}_{q}\left[\sum\nolimits_{t=2}^{T}\mathrm{D}_{\mathrm{KL}}(q({x}_{t-1}\mid{x}_{t},{x}_{0})\parallel p_{\theta}({x}_{t-1}\mid{x}_{t},t))\right]$
@@
To better control noise addition, GNDiff employs a noise schedule (He et al. 2022) that measures noise at each step by masking and evenly distributing information. Tokens are assigned varying mask probabilities, ordered by informativeness into $T$ stages. The forward process begins by masking the most informative tokens and ends with the least informative, simplifying the reverse process. Specifically, mask information is evenly distributed during the forward step, as shown in Equation 9:
$\sum\nolimits_{i=1}^{n}\overline{\alpha}_{t}^{i}H(x_{0}^{i})=\left(1-\frac{t}{T}\right)\sum\nolimits_{i=1}^{n}H(x_{0}^{i}),$ (9)
where $H$ represents entropy, quantifying the information content of a node. $x_{0}^{i}$ denotes the $i$th token in the sequence and $n$ is the sequence length. $\overline{\alpha}_{t}^{i}=\prod_{j=1}^{t}(1-\beta_{j}^{i})$ is the probability that the node retains its original state at time step $t$, i.e. $x_{t}^{i}=x_{0}^{i}$. We expect that if $H(x_{t}^{i})<H(x_{t}^{j})$, then $\bar{\alpha}_{t}^{i}>\bar{\alpha}_{t}^{j}$, so low-information tokens appear earlier in the reverse process than high-information ones. In practice, a token’s entropy $H(x)$ is computed as the negative logarithm of its frequency in the training corpus. Based on these properties, we construct $\overline{\alpha}_{t}^{i}$ as follows:
$\overline{\alpha}_{t}^{i}=1-\frac{t}{T}-G(t)\cdot\tilde{H}({x}_{0}^{i}),$ (10)
$G(t)=\mu\sin\frac{t\pi}{T},$ (11)
$\tilde{H}({x}_{0}^{i})=1-\frac{\sum_{j=1}^{n}H({x}_{0}^{j})}{nH({x}_{0}^{i})},$ (12)
where $G(t)$ controls the influence of information at time step $t$, with its effect modulated by the hyperparameter $\mu$.
Dual-Domain Periodic Contrastive Learning
Inspired by Xu et al. (2023), we propose a dual-domain periodic contrastive learning method. Unlike the original approach, we use a dual-domain mapping strategy because we observed that specific queries involve multiple periodic events that are challenging to distinguish. To address this issue, we map entities from periodic and non-periodic events to Poincaré and Euclidean spaces, respectively, to identify highly correlated entities within each domain. By leveraging the spatial characteristics of Poincaré space, we more accurately capture the relationships between periodic entities and distinguish similar ones.
@@
$\mathbf{d}_{p}(s,o)=cosh^{-1}\Bigg{(}1+2\frac{\parallel s-o\parallel_{2}^{2}}{\left(1-\parallel s\parallel_{2}^{2}\right)\left(1-\parallel o\parallel_{2}^{2}\right)}\Bigg{)},$ (15)
@@
$\mathbf{S}_{np}^{s,r}(o_{i})=tanh(\mathbf{W}_{np}{s}\oplus r+\mathbf{b}_{np})\mathbf{E}^{T}-\mathbf{Z}_{t}^{s,r}+d_{{E}}(s,o_{i}).$ (17)
$\mathcal{L}_{\mathrm{ce}}=-\sum_{q\in I}\log\{\frac{exp(S_{p}^{s,r}(o_{i}))}{\sum\limits_{o_{j}\in\mathcal{E}}exp(S_{p}^{s,r}(o_{j}))}+\frac{exp(S_{np}^{s,r}(o_{i}))}{\sum\limits_{o_{j}\in\mathcal{E}}exp(S_{np}^{s,r}(o_{j}))}\}.$ (18)
Here, $o_{i}$ denotes the ground truth object entity for a given query $q$. The goal of $\mathcal{L}_{ce}$ is to distinguish the ground truth from other values by comparing each scalar in $S_{p}^{s,r}$ and $S_{np}^{s,r}$.
$\mathcal{L}_{sup}=\sum_{q\in I}\frac{-1}{\mid P(i)\mid}\sum_{p\in P(i)}\log\frac{\exp(z_{q}\cdot z_{p}/\tau)}{\sum\exp(z_{q}\cdot z_{a}/\tau)},$ (19)
where $I$ is the set of positive samples, $P(i)$ is the set of queries with the same label in the small batch, $z_{q}$ denotes the embedding vector of query $q$, $z_{p}$ represents the embedding vector of queries sharing the same label as $q$, $z_{a}$ denotes the embedding vector of all non-$q$ queries in the small batch.
Method ICEWS18 ICEWS14 WIKI
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
ComplEx(2016) 30.09 21.88 34.15 45.96 24.47 16.13 27.49 41.09 47.84 38.15 50.08 58.46
R-GCN(2018) 23.19 16.36 25.34 36.48 26.31 18.23 30.43 45.34 37.57 28.15 39.66 ——
ConvE(2018) 36.67 28.51 39.80 50.69 40.73 33.20 43.92 54.35 47.57 38.76 50.1 63.74
RE-Net(2019) 42.93 36.19 45.47 55.80 45.71 38.42 49.06 59.12 51.97 48.01 52.07 53.91
xERTE(2020) 36.95 30.71 40.38 49.76 32.92 26.44 36.58 46.05 58.75 58.46 58.85 59.34
CyGNet(2021) 46.69 40.58 49.82 57.14 48.63 41.77 52.50 60.29 45.5 50.48 50.79 68.95
EvoKG(2022) 29.67 12.92 33.08 58.32 18.30 6.30 19.43 39.37 50.66 12.21 63.84 64.73
@@
DPCL-Diff 47.02 40.03 49.93 60.66 66.59 62.89 67.79 73.64 68.44 68.41 68.45 68.55
Improve(%) 7.55 $\uparrow$ 5.59 $\uparrow$ 9.38$\uparrow$ 10.71$\uparrow$ 28.23 $\uparrow$ 30.37 $\uparrow$ 28.37 $\uparrow$ 24.12 $\uparrow$ 2.35 $\uparrow$ 2.46 $\uparrow$ 2.26 $\uparrow$ 2.34 $\uparrow$
Table 1: Experimental results of temporal link prediction on three datasets TKGs. The time-filtered MRR, H@1, H@3, and H@10 metrics are multiplied by 100. The best results are in bold, and the previous best results are underlined. Improve refers to the improvement of DPCL-Diff compared to our baseline CENET.
Training Details and Inference
@@
$\mathcal{L}_{dpcl}=\alpha\cdot\mathcal{L}_{diff}+(1-\alpha)\cdot(\mathcal{L}_{ce}+\mathcal{L}_{sup}),$ (20)
@@
$\mathbf{P}_{\text{diff}}(o_{i}\mid q)=\prod\nolimits_{t=1}^{T}p_{\theta}(x_{t-1}\mid x_{t}),$ (21)
@@
$\mathbf{P}_{dpcl}(o_{i}\mid q)=\frac{\exp(s(o_{i}\mid q))}{\sum_{o_{j}\in\mathcal{E}}\exp(s(o_{j}\mid q))}.$ (22)
@@
Experiments
Experimental Setup
Datasets and Evaluation Metrics. To evaluate the effectiveness of the proposed DPCL-Diff, we report the reasoning performance on four public datasets: ICEWS14 (Trivedi et al. 2017), ICEWS18 (Jin et al. 2019), WIKI (Leblay and Chekol 2018), and YAGO (Mahdisoltani, Biega, and Suchanek 2013). We divided all datasets into three groups, roughly partitioning them into train (80%), validation (10%), and test (10%) sets by timestamp. The partitioning of the ICEWS14 dataset differs slightly from the original dataset. Table 2 shows more detailed information about the datasets. In our experiments, we chose the mean reversed rank (MRR) and the hit rate @1,3,10 (Liang et al. 2023) as the metrics.